Draw a Lewis structure for C2H3Cl. Show all unshar
Answer #1
Concepts and reason

A Lewis structure is a structural representation of the molecule including the electron distribution around the atoms. The valence electrons on each atom can be found using the atomic number of each atom.

Fundamentals

Lewis structure includes all non-bonded electrons in the structure. Each atom in the molecule should fulfill the octet rule.

The dots are used to represent the lone pairs or non-bonded pairs of electrons and the dot pairs or a line is used to represent the covalent bond between two atoms. Nonbonding electrons should exist as pairs.

The valance electron of each atom is required for drawing the Lewis structure.

The given molecule is C2H3Cl{{\rm{C}}_{\rm{2}}}{{\rm{H}}_{\rm{3}}}{\rm{Cl}}

Totalnumberofvalenceelectrons=2(4)+3(1)+1(7)=18\begin{array}{l}\\{\rm{Total}}\,{\rm{number}}\,{\rm{of}}\,{\rm{valence}}\,{\rm{electrons}}\,{\rm{ = }}\,{\rm{2}}\left( {\rm{4}} \right){\rm{ + 3}}\left( {\rm{1}} \right){\rm{ + 1}}\left( {\rm{7}} \right)\\\\\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\rm{ = 18}}\\\end{array}

Total number of valance electrons of C2H3Cl{{\rm{C}}_{\rm{2}}}{{\rm{H}}_{\rm{3}}}{\rm{Cl}} =18

The skeleton structure for the given molecule is as follows:

Figure 1

The molecule showing all nonbonding electrons on the respective atoms is as follows:

I
H
Figure 2

The Lewis structure is as follows:

10:
Lewis structure
Figure 2

Ans:

Lewis structure of C2H3Cl{{\rm{C}}_{\rm{2}}}{{\rm{H}}_{\rm{3}}}{\rm{Cl}} is as follows:

工、